# RBLA: Rank-Based-LoRA-Aggregation for Fine-tuning Heterogeneous Models in FLaaS

### v1

@@ 5 Experiments and Evaluation
In our study, we evaluate the effectiveness of RBLA using MLP and CNN architectures on the MNIST, FMNIST, CIFAR-10 and CINIC-10 [4] datasets, comparing its performance against two baselines: Zero-Padding (ZP) [2] and Full Fine-Tune (FFT) FedAvg [22]. The models are implemented using the TensorFlow library in Python, with three neural network models included. Additionally, all experiments are conducted with a fixed seed of 42 to ensure reproducibility. All experiments Within the same subfigure share the same configuration, differing only in the aggregation algorithms used.
@@ 5.1 Experiments setup
In the experiment, the MNIST MLP model consists of two hidden layers with 200 neurons each, activated by ReLU, and a 10-class softmax output layer suitable for flattened 28x28 pixel images (784-dimensional vectors). The MNIST CNN model comprises two convolutional layers with 32 and 64 filters, followed by MaxPooling layers, a 512-unit fully connected layer, and a 10-class softmax output layer designed for 28x28 inputs. Both models were trained with a batch size 64 using the SGD optimizer with a learning rate of 0.01. For the CIFAR dataset, we used a model that includes three convolutional layers: batch normalization, max pooling, dropout, and fully connected layers. The first two sets of convolutional layers have 32 and 64 filters with a 3x3 kernel size and ReLU activation, followed by pooling layers for downsampling and dropout layers for regularization. The feature maps are then flattened and passed through two fully connected layers with 512 neurons each, followed by dropout, and finally output through a 10-class softmax layer. The model we used for CINIC dataset has two extra dense layers with 512 neurons each compared with CIFAR model. The optimizer for the CIFAR and CINIC experiments was set to Adam [14], and LoRA is only applied on dense layers for all experiments. Additionally, we tested two participation scenarios: one where all clients participate in each training round and another where 20% of clients are randomly selected to participate in each communication round.
@@ 5.3 Evaluation Results
In this section, we present the evaluation results across various datasets with different configurations. Table. 1 shows the exact number of communication rounds taken for each method to reach the target test accuracy of the global model. Figure. 5 to Figure. 10 shows the learning curve of the global model’s test accuracy to training rounds. As an example, the target test accuracy set for the MNIST MLP experiment is 95%, the maximum test accuracy of ZP can reach in the training is 94.87%, which presented in the parentheses, while FFT reaches 95.04% test accuracy in 40th training rounds and RBLA reaches 95.06% in 11th training rounds. To ensure clear visual comparisons, a rolling average with a window size of 10 was applied to smooth the data, represented by solid lines, and the dotted lines illustrate the original and unsmoothed results. The effectiveness of RBLA is assessed on the MNIST, FMNIST, and CIFAR-10 datasets using MLP and CNN models under both full participation and random selection settings. These experiments compare the performance of RBLA against zero-padding and full model fine-tuning.
Table 1: The minimum number of training rounds used for each method to achieve the target test accuracy for the global model. N/A indicates that the corresponding aggregation strategy cannot achieve the target accuracy in 50 training rounds. All experiments in the same column are conducted under the same setting.
@@
In the MNIST MLP experiments, RBLA demonstrates a superior convergence rate under the full participate setting, reaching a test accuracy of 95% at the 11th communication round. In contrast, zero-padding and FFT require 42 and 40 rounds, respectively, to reach similar accuracy levels. This indicates that RBLA reduces the required training rounds by 40% compared to zero-padding. In the random selection scenario, RBLA maintains a higher accuracy with less fluctuation throughout the training process compared to zero-padding and achieves a performance close to FFT, which exhibits better variability.
@@
Figure 5: Evaluation of RBLA for MNIST dataset with MLP model where a) all clients participate in training, b) a random set of clients are selected for training.
@@
Figure 6: Evaluation of RBLA for FMNIST dataset with MLP model where a) all clients participate in training, b) a random set of clients are selected for training.
(a) All participate
@@
Figure 7: Evaluation of RBLA for MNIST dataset with CNN model where a) all clients participate in training, b) a random set of clients are selected for training.
@@
Figure 8: Evaluation of RBLA for FMNIST dataset with CNN model where a) all clients participate in training, b) a random set of clients are selected for training.
Figure. 6 shows the FMNIST MLP experiments, RBLA again outperforms the other methods, reaching 80% test accuracy by the 7th communication round, whereas FFT achieves the same accuracy after 4 rounds, and zero-padding fails to reach this target accuracy. This demonstrates that RBLA can reduce the number of training rounds by 36% compared to FFT and performs significantly better than zero-padding. Figures. 7 and 8 show the experimental results on CNN models using MNIST and FMNIST, respectively. In both cases, the advantage of RBLA becomes even more evident. In the MNIST full participation setting, RBLA achieves 98% test accuracy by the 4th communication round, which is faster than FFT by 18 rounds and faster than zero-padding by 7 rounds. Under the random selection setting, RBLA’s learning curve closely matches that of FFT and outperforms zero-padding.
@@
Figure 9: Evaluation of RBLA for CINIC-10 dataset with CNN model where a) all clients participate in training, b) a random set of clients are selected for training.
@@
Figure 10: Evaluation of RBLA for CIFAR-10 dataset with CNN model where a) all clients participate in training, b) a random set of clients are selected for training.
In the FMNIST CNN experiments, as shown in Figure. 8, RBLA shows its efficiency by achieving 98% test accuracy of the global model at the 7th communication round, outperforming zero-padding, which only reaches 97% test accuracy at the 24th communication round, while FFT fails to converge to the target accuracy. These results consistently demonstrate that RBLA converges faster and maintains higher stability across different datasets and model types. The significant reduction in training rounds and the lower fluctuation in test accuracy further confirm the effectiveness of RBLA over traditional zero-padding methods, particularly in non-IID settings. Finally, the CIFAR-10 and CINIC-10 experiments are shown in Figure. 10 and Figure. 9, respectively. RBLA gradually closes the performance gap with FFT as training rounds increase, outperforming zero-padding, while RBLA and FFT show similar convergence speeds at the beginning of training, RBLA converges more slowly after several rounds in CIFAR-10 experiment. But exhibit a faster convergence rate than both zero-padding and FFT in CINIC-10 experiments.

### v2

@@ 5 Experiments and Evaluation
In our study, we evaluate the effectiveness of RBLA using MLP and CNN architectures on the MNIST, FMNIST, CIFAR-10, and CINIC-10 [5] datasets, comparing its performance against two baselines: Zero-Padding (ZP) [3] and Full Fine-Tune (FFT) FedAvg [24]. The models are implemented using the TensorFlow library in Python, with three neural network models included. Additionally, all experiments are conducted with a fixed seed of 42 to ensure reproducibility. All experiments within the same subfigure share the same configuration, differing only in the aggregation algorithms used.
@@ 5.1 Experiments setup
In the experiment, the MNIST MLP model consists of two hidden layers with 200 neurons each, activated by ReLU, and a 10-class softmax output layer suitable for flattened 28x28 pixel images (784-dimensional vectors). The MNIST CNN model comprises two convolutional layers with 32 and 64 filters, followed by MaxPooling layers, a 512-unit fully connected layer, and a 10-class softmax output layer designed for 28x28 inputs. Both models were trained with a batch size of 64 using the SGD optimizer with a learning rate of 0.01. For the CIFAR dataset, we used a model that includes three convolutional layers: batch normalization, max pooling, dropout, and fully connected layers. The first two sets of convolutional layers have 32 and 64 filters with a 3x3 kernel size and ReLU activation, followed by pooling layers for downsampling and dropout layers for regularization. The feature maps are then flattened and passed through two fully connected layers with 512 neurons each, followed by dropout, and finally output through a 10-class softmax layer. The model that we used for the CINIC dataset has two extra dense layers with 512 neurons each compared to the CIFAR model. The optimizer for the CIFAR and CINIC experiments was set to Adam [16], and LoRA is applied only to dense layers for all experiments. Additionally, we tested two participation scenarios: one where all clients participate in each training round and another where 20% of clients are randomly selected to participate in each communication round.
@@ 5.3 Evaluation Results
Table 1: The minimum number of training rounds used for each method to achieve the target test accuracy for the global model in full participation experiments. N/A indicates that the corresponding aggregation strategy cannot achieve the target accuracy within 50 training rounds. All experiments in the same column are conducted under the same settings.
@@
In this section, we present the evaluation results across various datasets with different configurations. Table. 1 shows the exact number of communication rounds taken for each method to reach the target test accuracy of the global model. The target accuracy was selected based on the global model accuracy at the communication round where RBLA’s convergence speed significantly decreases, and this accuracy can also be reached or nearly reached by the other two methods. Figure. 5 to Figure. 10 show the learning curve of the global model’s test accuracy to training rounds; the left sub-figure shows the result with full participants, and the right sub-figure shows the result with random client selection. As an example, the target test accuracy set for the MNIST MLP experiment is 95%, where the convergence of RBLA significantly reduced. The maximum test accuracy of ZP can reach during training is 94.87% (presented in the parentheses), while FFT reaches 95.04% test accuracy in the 40th training round, and RBLA reaches 95.06% in the 11th training round. To ensure clear visual comparisons, a rolling average with a window size of 10 was applied to smooth the data, represented by solid lines, and the dotted lines illustrate the original and unsmoothed results. The effectiveness of RBLA is assessed on the MNIST, FMNIST, and CIFAR-10 datasets using MLP and CNN models under both full participation and random selection settings. These experiments compare the performance of RBLA against zero-padding and full model fine-tuning.
@@
Figure 5: Evaluation of RBLA for MNIST dataset with MLP model.
@@
Figure 6: Evaluation of RBLA for FMNIST dataset with MLP model.
(a) Full participate
@@
Figure 7: Evaluation of RBLA for MNIST dataset with CNN model.
@@
Figure 8: Evaluation of RBLA for FMNIST dataset with CNN model.
In the MNIST MLP experiments, RBLA demonstrates a superior convergence rate under the full participation setting, reaching a test accuracy of 95% at the 11th communication round. In contrast, zero-padding and FFT require 42 and 40 rounds, respectively, to reach similar accuracy levels. This indicates that RBLA reduces the required training rounds by 40% compared to zero-padding. In the random selection scenario, RBLA maintains a higher accuracy with less fluctuation throughout the training process compared to zero-padding and achieves performance close to FFT, which exhibits better variability.
@@
Figure 9: Evaluation of RBLA for CINIC-10 dataset with CNN model.
@@
Figure 10: Evaluation of RBLA for CIFAR-10 dataset with CNN model.
Figure. 6 shows the FMNIST MLP experiments, where RBLA again outperforms the other methods, reaching 80% test accuracy by the 7th communication round, whereas FFT achieves the same accuracy after 4 rounds, and zero-padding fails to reach this target accuracy. This demonstrates that RBLA can reduce the number of training rounds by 36% compared to FFT and performs significantly better than zero-padding. Figures. 7 and 8 show the experimental results on CNN models using MNIST and FMNIST, respectively. In both cases, the advantage of RBLA becomes even more evident. In the MNIST full participation setting, RBLA achieves 98% test accuracy by the 4th communication round, which is faster than FFT by 18 rounds and faster than zero-padding by 7 rounds. Under the random selection setting, RBLA’s learning curve closely matches that of FFT and outperforms zero-padding. In the FMNIST CNN experiments, as shown in Figure. 8, RBLA shows its efficiency by achieving 98% test accuracy of the global model at the 7th communication round, outperforming zero-padding, which only reaches 97% test accuracy at the 24th communication round, while FFT fails to converge to the target accuracy. These results consistently demonstrate that RBLA converges faster and maintains higher stability across different datasets and model types. The significant reduction in training rounds and the lower fluctuation in test accuracy further confirm the effectiveness of RBLA over traditional zero-padding methods, particularly in non-IID settings. Finally, the CIFAR-10 and CINIC-10 experiments are shown in Figures 10 and 9, respectively. RBLA gradually closes the performance gap with FFT as training rounds increase, outperforming zero-padding. While RBLA and FFT show similar convergence speeds at the beginning of training, RBLA converges more slowly after several rounds in the CIFAR-10 experiment but exhibits a faster convergence rate than both zero-padding and FFT in the CINIC-10 experiments.